OpenStudy (anonymous):

how to find the derivative of radical (1+e^2x / 1-e^2x)??

OpenStudy (anonymous):

is the whole thing in the radical?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{\frac{1+e^{2x}}{1-e^{2x}}}\]??

OpenStudy (anonymous):

^^ Yeah

OpenStudy (anonymous):

yeaaa

OpenStudy (anonymous):

^^ Agreed lol